LOIRP Status Report 5 September 2013

Dennis Wingo: Status end of day, September 5, 2013
Got through image G5-071 today. The head held out and maybe we have a few more left in old bessie…..
We went to see our Copernicus image, the big 23 gigabyte one, on the NASA Hyperwall 2 today. Got some good pictures, Keith will post them tomorrow.
Will post the image numbers tomorrow as well….
– Status mid day, September 5, 2013
A funny thing happened on the way to the tapes, we had a whole tape with nothing but black, an unexposed image. This would have been LOV-218H.
The next image, 216M has some weird pattern on the image, which according to the ground team chatter is the Bimat supply roll. Apparently this the last two images wound around the supply roll which cased the weird pattern. You can see this pattern on the existing image at: http://www.lpi.usra.edu/resources/lunarorbiter/images/preview/5216_med.jpg
This means that we are half way through the LOV film roll and are now running the final readout images. Today we were able to skip G5-060-63 because 60-62 were duplicates with Woomera images and G5-063 had nothing but a blank image on it. Thus our first real capture today is on G5-064. As stated yesterday, the halfway point in tape count is G5-079 so we obviously have many tapes and overlaps to go. This makes tape 227 out of 478 for LOV, so we are making progress! The head is not going to last much longer though, possibly no longer than today.